Find the moment coefficient of Skewness and kurtosis for the dat below. a) 9, 3, 4, 2, 9, 5,



1
Expert's answer
2022-03-14T17:27:12-0400

Given.

9,3,4,2,9,5


(a) Skewness


"\mu_3=\dfrac{\sum (x_i-\mu)^3}{(n-1)\sigma^3}"


"\mu = \dfrac{\sum X_i}{N}"


"\mu=\dfrac{9+3+4+2+9+5}{6} =5.33"


"\sigma=\sqrt\dfrac{(X_i -u)^2}{N}"


"\sigma=\sqrt \dfrac{45.34}{6}= 2.75"


"\mu_3=\dfrac{46.57}{5(2.75)^3}=0.448"


(b) Kurtosis


"Kurt (X)=\dfrac{\mu^4}{\sigma^4}"


"\mu^4=\dfrac{\sum(X_i -\mu)^4}{N}"


"\mu^4=\dfrac{518.39}{6}=86.40"


"Kurt (X) =\dfrac{86.40}{57.19}=1.51"